OpenStudy (anonymous):

Is this correct so far? The Nut Barn requires 15 pounds of mixed walnuts and pecans from their wholesaler. Walnuts cost $2.26 per pound and pecans cost $1.96 per pound. How many pounds of pecans are needed if they only want to pay $2.18 per pound for the mixture?

OpenStudy (anonymous):

p = pecans w = walnuts Two Equations: w + p = 15 2.26w + 1.96p = 2.18(w + p) Solve for w in first equation. w + p = 15 subtract ----------- w = p - 15 Plug in this new information into second equation. 2.26(p - 15) + 1.96p = 2.18((p - 15) + p) distribute ----------------------------------------- 2.26p - 33.9 + 1.96p = 2.18((p - 15) + p) subtract ----------------------------------------- 2.26p + 1.96p = 2.18((p - 15) + p) + 33.9

OpenStudy (phi):

it looks ok, but I would replace (w+p) in the 2nd equation with 15 immediately (you know w+p=15)

OpenStudy (anonymous):

2.26(p - 15) + 1.96p = 2.18(15) multiply ------------------------------- 2.26(p - 15) + 1.96p = 37.2 distribute ------------------------------- 2.26p - 33.9 + 1.96p = 37.2 subtract ------------------------------- 2.26p + 1.96p = -71.1 add ------------------------------- 4.22p = -71.1 divide ------------------------------- p = 16.84834123

OpenStudy (anonymous):

I did that and solved, but got a bunch of decimal points... What's wrong?

OpenStudy (phi):

how did you get -71.1?

OpenStudy (anonymous):

I subtracted -33.9 on both sides

OpenStudy (phi):

I think you should add +33.9 to both sides that way the -33.9 + 33.9 on the left will add to zero.

OpenStudy (anonymous):

ahhhh Right. Okay I'll do that now. :)

OpenStudy (phi):

no, wait... first fix w+p= 15 add -p to both sides w+p-p= 15-p or w= 15-p you are using w= p-15 which is not good.

OpenStudy (anonymous):

Got it. P = 11. :) Thanks!